Find the centre and radius of circle whose equation is given by 2x*2+2y*2-6x=36
Answer by Alan3354(69443)

Find the centre and radius of circle whose equation is given by 2x*2+2y*2-6x=36
--------
x^2 + y^2 - 3x = 18
x^2-3x + y^2 = 18
Complete the squares for x and y
x^2+3x+2.25 + y^2 = 20.25 = 4.5^2
(x+1.5)^2 + y^2 = 4.5^2
Center (-1.5,0)
r = 4.5
==========
Use ^ for exponents, * for multiply